OpenStudy (anonymous):

Factor the trinomial 3x^2 + 17x - 6

pooja195 (pooja195):

3x^2 - 17x - 6 = ( 3x + 1)( x - 6)

pooja195 (pooja195):

im hoping uvkno the step??

OpenStudy (anonymous):

Not really....

pooja195 (pooja195):

3x stays the same u kno that rite

pooja195 (pooja195):

3-2 is 1 u add

pooja195 (pooja195):

cancel out the 17

pooja195 (pooja195):

and the that leave u with x-6

OpenStudy (anonymous):

how do you know to subtract 3-2?

OpenStudy (tkhunny):

Factors of 3 1, 3 Factors of -6 1, -6 -1, 6 2,-3 -2,3 Which combination produces +17 in the middle? 3*3 = 9 1*2 = 2 Okay, all the 2-3s are out. We're down to this: (x + 1)(3x - 6) = ?? (x - 1)(3x + 6) = ?? Which one works?

OpenStudy (anonymous):

Good lord, I am lost.......

OpenStudy (tkhunny):

Just enumerate the possibilities. In this case, there are only four. Simply try them out!